1. Sketch the curve given by the parametric equations:

  \begin{cases} & x= 2cos(t) \\ & y= 3sin(t) \\ & 0\leq x \leq \frac{\pi }{4} \end{cases}(please use dashed lines on graph f or t beyond 4 )

by eliminating the parameter ?. Label the direction a point on the curve moves as ? increases.

PLEASE GIVE FULL WORKING FOR THE QUESTION.
Answer #1

SOLUTION:-

Given :-

\left\{\begin{matrix} x=2\cos{t}\\ y=3\sin{t}\\ 0\leq t\leq \frac{ \Pi}{4} \end{matrix}\right.

Consider,

\\ \frac{x^{2}}{4}+\frac{y^{2}}{9}=\frac{4\cos^{2}{t}}{4}+\frac{9\sin^{2}{t}}{9}

\\ \frac{x^{2}}{4}+\frac{y^{2}}{9}= \cos^{2}{t}+\sin^{2}{t}=1

\\ \frac{x^{2}}{4}+\frac{y^{2}}{9}=1

Which is an equation of an ellipse having major axis along y -axis of 3 units and minor axis along x-axis of 2 units.

\\ at\; t=0\\ x=2\cos{0}=2\\ y=2\sin{0}=0

\\ at\; t=\frac{\pi}{4}\\ x=2\cos{\frac{\pi}{4}}=\sqrt{2}\\\\ y=2\sin{\frac{\pi}{4}}=\frac{3}{\sqrt{2}}

THE REQUIRED GRAPH IS :-

t = 니 153름 2 (20) 3 키 - 3 진 . - 2 3 CS Scanned with CamScanner